OpenStudy (anonymous):

f(x)=2-x and g(x)=2x^2+x+9 (f(g))(x) , (g(f))(x) and (f(g))(3)

OpenStudy (anonymous):

hey aditya can you help?

OpenStudy (anonymous):

(f(g))(x)=2-(2x^2+x+9) =2-2x^2-x-9 =-2x^2-x-7 (f(g))(3)=-2*3*3-3-7 =-18-10 =-28 (g(f))(x)=2(2-x)^2+(2-x)-7 =2(4-4x+x^2)+2-x-7 =8-8x+2x^2-x-5 =2x^2-9x+3